OpenStudy (anonymous):

Solve The Trinomial equation k^2+8k=84

OpenStudy (callisto):

k^2+8k=84 k^2+8k -84=0 (k-6)(k+14) =0 (k-6)=0 or (k+14) =0 Can you solve them from here the get the answers?

OpenStudy (anonymous):

Yes I can

OpenStudy (callisto):

Good :)